# Local geometry of feasible regions via smooth paths

### v1

@@ 2.1 Euclidean spaces
$\langle v,x-\bar{x}\rangle\leq o(\|x-\bar{x}\|)\quad\text{for all}~{}x\in C.$
@@ 2.2 Metric spaces
$d\big{(}\gamma(t_{1}),\gamma(t_{2})\big{)}=|t_{1}-t_{2}|\quad\text{for all}~{}t_{1},t_{2}\in[0,L(\gamma)].$
@@
$d\big{(}\gamma(t_{1}),\gamma(t_{2})\big{)}=|t_{1}-t_{2}|\cdot d(x,y)\quad\text{for all}~{}t_{1},t_{2}\in[0,1].$
@@ 3 Smooth approximate convexity in Euclidean spaces
Based on these ideas, we introduce a new definition that not only captures approximate convexity but in a $\mathcal{C}^{(1)}$-smooth sense. Given a set $C\subseteq{\mathbf{R}}^{n}$, as in [19], we call a curve $\gamma:[a,b]\rightarrow C$ a smooth curve segment in $C$ if it has an extension to a $\mathcal{C}^{(1)}$-smooth map from a neighborhood of $[a,b]$ to ${\mathbf{R}}^{n}$.
@@
Proof  Let $\bar{y}=H(\bar{x})$ and $M=H(U)$. Suppose $H=(h_{1},\dots,h_{m})$. Fix any $\epsilon^{\prime}>0$. By the injectivity of $\nabla H(\bar{x})$ and $\mathcal{C}^{(1)}$-smoothness of $H$, we find a neighborhood $\widetilde{U}\subseteq U$ of $\bar{x}$ in ${\mathbf{R}}^{n}$ and constants $c_{1},c_{2}>0$ such that for all $x,x^{\prime}\in\widetilde{U}$,
@@
Let $\epsilon=\min\{\frac{c_{1}}{2c_{2}}\epsilon^{\prime},\frac{1}{2}\}$. Since $C$ is smoothly approximately convex at $\bar{x}$, there exists $\delta>0$ such that $B_{2\delta}(\bar{x})\subseteq\widetilde{U}$ and for all $x,x^{\prime}\in B_{\delta}(\bar{x})\cap C$, there is an $\epsilon$-path from $x$ to $x^{\prime}$. Since $H:U\rightarrow M$ is a homeomorphism, there exists $\delta^{\prime}>0$ such that
$H^{-1}(B_{\delta^{\prime}}(\bar{y})\cap M)\subseteq B_{\delta}(\bar{x}).$
@@ 3.1 Other regularity conditions
Applying [27, Theorem 2.39], $\mathrm{Range}(\nabla F(\bar{x}))$ and $\mathrm{int}(D)-F(\bar{x})$ can be separated by a hyperplane, namely there exist $a\in{\mathbf{R}}^{m}\backslash\{0\}$ and $\alpha\in{\mathbf{R}}$ such that
@@
Moreover, it is easy to deduce from (3.12) that $(y,z)\in N_{D}(F(\bar{x}))$. So by (3.9) we have $y=0$. Thus, $\widetilde{C}$ is amenable at 00. Applying Proposition 3.12, we know that $\widetilde{C}$ is smoothly approximately convex at 00. Consequently, by (3.11) and Lemma 3.5, $C$ is smoothly approximately convex at $\bar{x}$. $\Box$
@@
Proof  By [21, Proposition 3.15 and Remark 3.2], there exist $\sigma,r_{0}>0$ such that for all $r\in(0,r_{0})$, the set $C\cap\overline{B}_{r}(\bar{x})$ is closed with
@@
Hence, by [14, Lemma 1], $\gamma$ is a smooth curve segment in $C$.
@@ 4.1 Embedded submanifolds
The generalization to embedded submanifolds is straightforward. In Euclidean spaces, smooth approximate convexity is characterized by comparing the derivatives of the paths with those of line segments. In embedded manifolds, we simply replace the line segments with averaging maps.
We consider a $\mathcal{C}^{(3)}$-smooth embedded submanifold $\mathcal{M}$ in ${\mathbf{R}}^{n}$ and a point $\bar{x}\in\mathcal{M}$. In this case, every pair of points in $\mathcal{M}$ near $\bar{x}$ are connected by a unique geodesic [21], therefore also by a unique averaging map. Moreover, since $\mathcal{M}$ is prox-regular at $\bar{x}$ (see [27]), we can apply Theorem 3.19 to see that all averaging maps of $\mathcal{M}$ near $\bar{x}$ are smooth curve segments. Note that $\mathcal{C}^{(3)}$-smoothness of the manifold is assumed for simplicity. Most of the results below hold under weaker smoothness conditions.
@@
The notion of prox-regularity in Riemannian manifolds was introduced in [15] with a different definition involving the exponential mapping. However, prox-regular sets in the sense of [15] also satisfy the projection property in Definition 4.3 (see [15, Theorem 3.12]). Note also that Definition 4.3 implies that $C$ is locally closed at $\bar{x}$.
@@
Similar correspondences hold for amenability and prox-regularity. The proof for the former involves illuminating results on the structure of amenable sets. The proof for the latter reveals the metric nature of prox-regularity. To avoid distracting readers from the main focus of this paper, we present their proofs in the appendices.
@@ 4.2 General manifolds
where $r=d_{\mathcal{M}}(x,x^{\prime})$ and $g:[0,1]\rightarrow\mathcal{M}$ is the averaging map from $x$ to $x^{\prime}$. We call such a map $\gamma$ a weak $\epsilon$-path from $x$ to $x^{\prime}$ in $\mathcal{M}$.
@@
Proof  Using the Nash embedding theorem [26], we know there is a $\mathcal{C}^{(3)}$ isometric embedding $H$ from $M$ to a space ${\mathbf{R}}^{n}$. Let $\bar{w}=H(\bar{x})$, ~$\widetilde{\mathcal{M}}=H(\mathcal{M})$ and $\widetilde{C}=H(C)$. Then $\widetilde{\mathcal{M}}$ is a $\mathcal{C}^{(3)}$-smooth embedded submanifold of ${\mathbf{R}}^{n}$. Let $H^{-1}$ denote the inverse of the mapping $H:\mathcal{M}\rightarrow\widetilde{\mathcal{M}}$. Suppose $C\cap V=\{x\in V:F(x)\in D\}$, where $V$ is a neighborhood of $\bar{x}$ in $\mathcal{M}$, $F:V\rightarrow{\mathbf{R}}^{m}$ is a $\mathcal{C}^{(1)}$-smooth map and $D\subseteq{\mathbf{R}}^{m}$ is a closed convex set satisfying (4.2). Define $\widetilde{V}=H(V)$ and $\widetilde{F}=F\circ H^{-1}:\widetilde{V}\rightarrow{\mathbf{R}}^{m}$. Note that $d\widetilde{F}(\bar{w})=dF(\bar{x})\circ dH^{-1}(\bar{w})$. So $\widetilde{V}$ is a neighborhood of $\bar{w}$ in $\widetilde{M}$ satisfying
@@ 5 First-order optimality condition on Riemannian manifolds
where $C$ is a subset of a Riemannian manifold $\mathcal{M}$, and $f:\mathcal{M}\rightarrow{\mathbf{R}}$ is a $\mathcal{C}^{(1)}$-smooth function. To discuss the first-order optimality condition, we need to introduce the following property for a tangent vector $v$ at $x\in C$, which is equivalent to the fact that $v$ belongs to the Bouligand tangent cone to $C$ at $x$ as given in [18, Definition 3.8].
@@
$vf\geq 0\quad\text{for all}\quad v\in T_{C}(\bar{x}).$ (5.2)
This condition must hold at any local minimizer of $f$ on $C$. The proof is elementary (cf. [18, Proposition 3.9]).
@@
Proof  By the Nash embedding theorem [26], we know there is a $\mathcal{C}^{(3)}$ isometric embedding $H$ from $\mathcal{M}$ to a space ${\mathbf{R}}^{n}$. Let $\bar{w}=H(\bar{x})$, ~$\widetilde{\mathcal{M}}=H(\mathcal{M})$ and $\widetilde{C}=H(C)$. Then $\widetilde{\mathcal{M}}$ is a $\mathcal{C}^{(3)}$-smooth embedded submanifold of ${\mathbf{R}}^{n}$. Let $H^{-1}$ denote the inverse of the mapping $H:\mathcal{M}\rightarrow\widetilde{M}$ and define $\tilde{f}=f\circ H^{-1}$. Using the same argument as in Theorem 4.12, we can see that $\widetilde{C}\subseteq\widetilde{\mathcal{M}}$ is smoothly weakly convex at $\bar{w}$. Suppose there exists a vector $v\in T_{C}(\bar{x})$ such that $vf<0$. It is easy to check that
@@
By the smooth weak convexity of $\widetilde{C}$ at $\bar{w}$ and considering that $t\mapsto g_{k}(r_{k}t)$ are averaging maps, we know there exist $k\in{\mathbb{N}}$ and a smooth curve segment $\gamma:[0,1]\rightarrow\widetilde{C}$ such that $\|g_{k}^{\prime}(0)-\tilde{v}\|<\epsilon$ and
@@
Feasible curves of linear decrease, like that guaranteed in Proposition 5.2, open up the possibility of constructing descent steps using a backtracking line search, like that in [5, Section 4.5].
@@ Appendix A Appendix: amenability on Riemannian manifolds
Proof  Suppose $\mathcal{M}=H(B_{\delta}(0))$ around $\bar{x}$, where $B_{\delta}(0)\subseteq{\mathbf{R}}^{k}$ and $H:B_{\delta}(0)\rightarrow{\mathbf{R}}^{n}$ is an embedding. Let $\widetilde{C}=H^{-1}(C)$. Then there exists an open neighborhood $W$ of 00 in ${\mathbf{R}}^{k}$ such that $\widetilde{C}\cap W=\{w\in W:\widetilde{F}(w)\in D\}$, where $\widetilde{F}:=F\circ H$. Note that $\nabla\widetilde{F}(0)=dF(\bar{x})\nabla H(0)$ and $T_{\mathcal{M}}(\bar{x})=\mathrm{Range}(\nabla H(0))$. Hence the equivalence follows directly from applying Lemma 3.11 to $(W,\widetilde{F},D)$. The local connectedness of $F^{-1}(\mathrm{int}(D))$ comes from from applying Remark 3.13 to $(W,\widetilde{F},D)$. $\Box$
@@
Proof  Suppose $\mathcal{N}_{1}=H(B_{\delta}(0))$ around $\bar{x}$, where $B_{\delta}(0)$ is an open ball in ${\mathbf{R}}^{k}$ and $H:B_{\delta}(0)\rightarrow{\mathbf{R}}^{n}$ is a $\mathcal{C}^{(1)}$-smooth embedding. Let $\widetilde{H}=\mathrm{P}_{\mathcal{N}_{2}}\circ H$. Then by [23, Lemma 2.1], $\widetilde{H}$ is $\mathcal{C}^{(1)}$-smooth around $\bar{x}$ with
@@ Appendix B Appendix: prox-regularity on Riemannian manifolds
In what follows, we include the proof of Lemma 4.6, which crucially relies on a characterization of prox-regularity by comparing the intrinsic distance with the Riemannian distance as below (see [13]).
@@
Proof  The reverse implication can be deduced from [21, Remark 3.2] and [24, Theorem 1.6]. The direct implication follows from Proposition B.6. $\Box$
@@
Before justifying why prox-regularity implies finite extrinsic curvature, we recall some properties of geodesic metric spaces and smooth Riemannian manifolds.
@@
For a $\mathcal{C}^{(3)}$-smooth Riemannian manifold $\mathcal{M}$ and a point $\bar{x}\in\mathcal{M}$, there exists a ball centered at $\bar{x}$ that is both a $\operatorname{CAT}(\kappa)$ and a $\operatorname{CBB}(\kappa^{\prime})$ space for some suitable $\kappa,\kappa^{\prime}\in{\mathbf{R}}$ (see, e.g., [6, Chapter II.1, Appendix], [7], [2]).
@@
For any points $x,y\in B_{R}(\bar{x})$, denote the geodesic in $M$ from $x$ to $y$ as $\gamma_{x,y}$ and its length as $l_{x,y}$. We recursively define a sequence of continuous curves $\gamma_{n}:[0,1]\rightarrow\mathcal{M}$, where $n\in{\mathbb{N}}\cup\{0\}$, as follows:
@@
and $S$ has positive reach (see [11]) if there exists $\delta>0$ such that $P_{S}(y)$ is a singleton for all $y\in\mathcal{M}$ with $d_{\mathcal{M}}(y,S)<\delta$.
@@
Proof  It follows directly from the definitions that $3\Rightarrow 2$ and $4\Rightarrow 1$. Moreover, by [21, Remark 3.2], we know $2\Rightarrow 3$; by [24, Theorem 1.6], we know $3\Rightarrow 4$. Proposition B.2 gives the equivalence between $1$ and $2$. We conclude that the four properties are equivalent. $\Box$

### v2

@@ 2.1 Euclidean spaces
$\langle v,x-\bar{x}\rangle\leq o(\|x-\bar{x}\|)\qquad\text{for all}~{}x\in C.$
@@ 2.2 Metric spaces
$d\big{(}\gamma(t_{1}),\gamma(t_{2})\big{)}=|t_{1}-t_{2}|\qquad\text{for all}~{}t_{1},t_{2}\in[0,L(\gamma)].$
@@
$d\big{(}\gamma(t_{1}),\gamma(t_{2})\big{)}=|t_{1}-t_{2}|\cdot d(x,y)\qquad\text{for all}~{}t_{1},t_{2}\in[0,1].$ (2.1)
@@ 3 Smooth approximate convexity in Euclidean spaces
Based on these ideas, we introduce a new definition that not only captures approximate convexity but in a $\mathcal{C}^{(1)}$-smooth sense. Given a set $C\subseteq{\mathbf{R}}^{n}$, as in [20], we call a curve $\gamma:[a,b]\rightarrow C$ a smooth curve segment in $C$ if it has an extension to a $\mathcal{C}^{(1)}$-smooth map from a neighborhood of $[a,b]$ to ${\mathbf{R}}^{n}$. (Note that the mapping $\eta$ used in the Introduction to define smooth approximate convexity is such an extension.)
@@
Proof  Let $\bar{y}=H(\bar{x})$ and $\mathcal{M}=H(U)$. Suppose $H=(h_{1},\dots,h_{m})$. Fix any $\epsilon^{\prime}>0$. By the injectivity of $\nabla H(\bar{x})$ and $\mathcal{C}^{(1)}$-smoothness of $H$, we find a neighborhood $\widetilde{U}\subseteq U$ of $\bar{x}$ in ${\mathbf{R}}^{n}$ and constants $c_{1},c_{2}>0$ such that for all $x,x^{\prime}\in\widetilde{U}$,
@@
Let $\epsilon=\min\{\frac{c_{1}}{2c_{2}}\epsilon^{\prime},\frac{1}{2}\}$. Since $C$ is smoothly approximately convex at $\bar{x}$, there exists $\delta>0$ such that $B_{2\delta}(\bar{x})\subseteq\widetilde{U}$ and for all $x,x^{\prime}\in B_{\delta}(\bar{x})\cap C$, there is an $\epsilon$-path from $x$ to $x^{\prime}$. Since $H:U\rightarrow\mathcal{M}$ is a homeomorphism, there exists $\delta^{\prime}>0$ such that
$H^{-1}(B_{\delta^{\prime}}(\bar{y})\cap\mathcal{M})\subseteq B_{\delta}(\bar{x}).$
@@ 3.1 Other regularity conditions
We conclude this section by mentioning another regularity property, called subsmoothness, which can be regarded as a translation of prox-regularity in a first-order context. Namely, a set $C\subseteq{\mathbf{R}}^{n}$ is “subsmooth” at a point $\bar{x}\in C$ if it is locally closed there and given any $\rho>0$, there exists $\epsilon>0$ such that $\langle v,x^{\prime}-x\rangle\leq\rho\|x^{\prime}-x\|$ for all $x,x^{\prime}\in C\cap B_{\epsilon}(\bar{x})$ and all $v\in N^{c}_{C}(x)\cap\overline{B}_{1}(0)$, where $N^{c}_{C}(x)$ is the closed convex hull of $N_{C}(x)$ called the Clarke normal cone to $C$ at $x$. Subsmoothness is strictly weaker than prox-regularity and strictly stronger than super-regularity, intrinsic approximate convexity, and Clarke regularity. Moreover, a function $f:{\mathbf{R}}^{n}\to{\mathbf{R}}$ that is locally Lipschitz continuous at $\bar{x}\in{\mathbf{R}}^{n}$ is approximately convex at $\bar{x}$ if and only if its epigraph is subsmooth at $(\bar{x},f(\bar{x}))$. We refer to the survey [31] for a comprehensive study of subsmooth sets.
@@
Applying [29, Theorem 2.39], $\mathrm{Range}(\nabla F(\bar{x}))$ and $\mathrm{int}(D)-F(\bar{x})$ can be separated by a hyperplane, namely there exist $a\in{\mathbf{R}}^{m}\backslash\{0\}$ and $\alpha\in{\mathbf{R}}$ such that
@@
Moreover, it is easy to deduce from (3.12) that $(y,z)\in N_{D}(F(\bar{x}))$. So by (3.9) we have $y=0$. Thus, $\widetilde{C}$ is amenable at 00. Applying Proposition 3.12, we know that $\widetilde{C}$ is smoothly approximately convex at 00. Consequently, by (3.11) and Lemma 3.5, $C$ is smoothly approximately convex at $\bar{x}$.
$\Box$
@@
Proof  By [22, Proposition 3.15 and Remark 3.2], there exist $\sigma,r_{0}>0$ such that for all $r\in(0,r_{0})$, the set $C\cap\overline{B}_{r}(\bar{x})$ is closed with
@@
Hence, by [15, Lemma 1], $\gamma$ is a smooth curve segment in $C$.
@@ 4.1 Embedded submanifolds
The generalization to embedded submanifolds is straightforward. In Euclidean spaces, smooth approximate convexity is characterized by comparing the derivatives of the paths with those of line segments. In embedded manifolds, we simply replace the line segments with averaging maps (see (2.1)).
We consider a $\mathcal{C}^{(3)}$-smooth embedded submanifold $\mathcal{M}$ in ${\mathbf{R}}^{n}$ and a point $\bar{x}\in\mathcal{M}$. In this case, every pair of points in $\mathcal{M}$ near $\bar{x}$ are connected by a unique geodesic, therefore also by a unique averaging map. Moreover, since $\mathcal{M}$ is prox-regular at $\bar{x}$ (see [29]), we can apply Theorem 3.19 to see that all averaging maps of $\mathcal{M}$ near $\bar{x}$ are smooth curve segments. Note that $\mathcal{C}^{(3)}$-smoothness of the manifold is assumed for simplicity. Most of the results below hold under weaker smoothness conditions.
@@
The notion of prox-regularity in Riemannian manifolds was introduced in [16] with a different definition involving the exponential mapping. However, prox-regular sets in the sense of [16] also satisfy the projection property in Definition 4.3 (see [16, Theorem 3.12]). Note also that Definition 4.3 implies that $C$ is locally closed at $\bar{x}$.
@@
Similar correspondences hold for amenability and prox-regularity.
@@
The proof for the former involves illuminating results on the structure of amenable sets. In connection to the latter, Bangert [2] studied sets that are prox-regular at all their points and showed, based on [17], that this notion does not depend on the considered Riemannian metric. Instead of building a proof relying on this result, we justify Lemma 4.6 as a consequence of the fact that prox-regularity at some point implies finite extrinsic curvature there. To avoid distracting readers from the main focus of the paper, we discuss these proofs in the appendices.
@@ 4.2 General manifolds
where $r=d_{\mathcal{M}}(x,x^{\prime})$ and $g:[0,1]\rightarrow\mathcal{M}$ is the averaging map from $x$ to $x^{\prime}$ (see (2.1)). We call such a map $\gamma$ a weak $\epsilon$-path from $x$ to $x^{\prime}$ in $\mathcal{M}$.
@@
Proof  Using the Nash embedding theorem [28], we know there is a $\mathcal{C}^{(3)}$ isometric embedding $H$ from $\mathcal{M}$ to a space ${\mathbf{R}}^{n}$. Let $\bar{w}=H(\bar{x})$, ~$\widetilde{\mathcal{M}}=H(\mathcal{M})$ and $\widetilde{C}=H(C)$. Then $\widetilde{\mathcal{M}}$ is a $\mathcal{C}^{(3)}$-smooth embedded submanifold of ${\mathbf{R}}^{n}$. Let $H^{-1}$ denote the inverse of the mapping $H:\mathcal{M}\rightarrow\widetilde{\mathcal{M}}$. Suppose $C\cap V=\{x\in V:F(x)\in D\}$, where $V$ is a neighborhood of $\bar{x}$ in $\mathcal{M}$, $F:V\rightarrow{\mathbf{R}}^{m}$ is a $\mathcal{C}^{(1)}$-smooth map and $D\subseteq{\mathbf{R}}^{m}$ is a closed convex set satisfying (4.2). Define $\widetilde{V}=H(V)$ and $\widetilde{F}=F\circ H^{-1}:\widetilde{V}\rightarrow{\mathbf{R}}^{m}$. Note that $d\widetilde{F}(\bar{w})=dF(\bar{x})\circ dH^{-1}(\bar{w})$. So $\widetilde{V}$ is a neighborhood of $\bar{w}$ in $\widetilde{M}$ satisfying
@@ 5 First-order optimality condition on Riemannian manifolds
where $C$ is a subset of a Riemannian manifold $\mathcal{M}$, and $f:\mathcal{M}\rightarrow{\mathbf{R}}$ is a $\mathcal{C}^{(1)}$-smooth function. To discuss the first-order optimality condition, we need to introduce the following property for a tangent vector $v$ at $x\in C$, which is equivalent to the fact that $v$ belongs to the Bouligand tangent cone to $C$ at $x$ as given in [19, Definition 3.8].
@@
$vf\geq 0\qquad\text{for all}~{}v\in T_{C}(\bar{x}).$ (5.2)
This condition must hold at any local minimizer of $f$ on $C$. The proof is elementary (cf. [19, Proposition 3.9]).
@@
Proof  By the Nash embedding theorem [28], we know there is a $\mathcal{C}^{(3)}$ isometric embedding $H$ from $\mathcal{M}$ to a space ${\mathbf{R}}^{n}$. Let $\bar{w}=H(\bar{x})$, ~$\widetilde{\mathcal{M}}=H(\mathcal{M})$ and $\widetilde{C}=H(C)$. Then $\widetilde{\mathcal{M}}$ is a $\mathcal{C}^{(3)}$-smooth embedded submanifold of ${\mathbf{R}}^{n}$. Let $H^{-1}$ denote the inverse of the mapping $H:\mathcal{M}\rightarrow\widetilde{\mathcal{M}}$ and define $\tilde{f}=f\circ H^{-1}$. Using the same argument as in Theorem 4.12, we can see that $\widetilde{C}\subseteq\widetilde{\mathcal{M}}$ is smoothly weakly convex at $\bar{w}$. Suppose there exists a vector $v\in T_{C}(\bar{x})$ such that $vf<0$. It is easy to check that
@@
By the smooth weak convexity of $\widetilde{C}$ at $\bar{w}$ and considering that $t\mapsto g_{k}(r_{k}t)$ are averaging maps (see (2.1)), we know there exist $k\in{\mathbb{N}}$ and a smooth curve segment $\gamma:[0,1]\rightarrow\widetilde{C}$ such that $\|g_{k}^{\prime}(0)-\tilde{v}\|<\epsilon$ and
@@
Feasible curves of linear decrease, like that guaranteed in Proposition 5.2, open up the possibility of constructing descent steps using a backtracking line search, like that in [6, Section 4.5].
@@ Appendix A Appendix: amenability on Riemannian manifolds
Proof  Suppose $\mathcal{M}=H(B_{\delta}(0))$ around $\bar{x}$, where $B_{\delta}(0)\subseteq{\mathbf{R}}^{k}$ and $H:B_{\delta}(0)\rightarrow{\mathbf{R}}^{n}$ is an embedding. Let $\widetilde{C}=H^{-1}(C)$. Then there exists an open neighborhood $W$ of 00 in ${\mathbf{R}}^{k}$ such that $\widetilde{C}\cap W=\{w\in W:\widetilde{F}(w)\in D\}$, where $\widetilde{F}:=F\circ H$. Note that $\nabla\widetilde{F}(0)=dF(\bar{x})\nabla H(0)$ and $T_{\mathcal{M}}(\bar{x})=\mathrm{Range}(\nabla H(0))$. Hence the equivalence follows directly from applying Lemma 3.11 to $(W,\widetilde{F},D)$. The local connectedness of $F^{-1}(\mathrm{int}(D))$ comes from applying Remark 3.13 to $(W,\widetilde{F},D)$.
$\Box$
@@
Proof  Suppose $\mathcal{N}_{1}=H(B_{\delta}(0))$ around $\bar{x}$, where $B_{\delta}(0)$ is an open ball in ${\mathbf{R}}^{k}$ and $H:B_{\delta}(0)\rightarrow{\mathbf{R}}^{n}$ is a $\mathcal{C}^{(1)}$-smooth embedding. Let $\widetilde{H}=\mathrm{P}_{\mathcal{N}_{2}}\circ H$. Then by [24, Lemma 2.1], $\widetilde{H}$ is $\mathcal{C}^{(1)}$-smooth around $\bar{x}$ with
@@ Appendix B Appendix: prox-regularity on Riemannian manifolds
The goal of this section is to discuss prox-regularity in connection to other geometric notions that measure how close a set is to being convex, which also allows us to derive the proof of Lemma 4.6. In particular, we focus on a characterization of prox-regularity by comparing the intrinsic distance with the Riemannian distance as below (see [14]).
@@
Proof  The reverse implication can be deduced from [22, Remark 3.2] and [25, Theorem 1.6]. The direct implication follows from Proposition B.6. $\Box$
@@
Although equivalences among related properties have been extensively studied in the literature in the setting of Euclidean spaces and Riemannian manifolds, we include here, based on similar techniques, a complete and direct proof of the fact that prox-regularity implies finite extrinsic curvature with the aim of clarifying details for the reader’s convenience. Before doing this, we recall some properties of geodesic metric spaces and smooth Riemannian manifolds.
@@
For a $\mathcal{C}^{(3)}$-smooth Riemannian manifold $\mathcal{M}$ and a point $\bar{x}\in\mathcal{M}$, there exists a ball centered at $\bar{x}$ that is both a $\operatorname{CAT}(\kappa)$ and a $\operatorname{CBB}(\kappa^{\prime})$ space for some suitable $\kappa,\kappa^{\prime}\in{\mathbf{R}}$ (see, e.g., [7, Chapter II.1, Appendix], [8], [3]).
@@
For any points $x,y\in B_{R}(\bar{x})$, denote the geodesic in $\mathcal{M}$ from $x$ to $y$ as $\gamma_{x,y}$ and its length as $l_{x,y}$. We recursively define a sequence of continuous curves $\gamma_{n}:[0,1]\rightarrow\mathcal{M}$, where $n\in{\mathbb{N}}\cup\{0\}$, as follows:
@@
and $S$ has positive reach (see [12]) if there exists $\delta>0$ such that $P_{S}(y)$ is a singleton for all $y\in\mathcal{M}$ with $d_{\mathcal{M}}(y,S)<\delta$. For sets of positive reach, a result related to Proposition B.6 is [26, Proposition 2.4].
@@
Proof  It follows directly from the definitions that $3\Rightarrow 2$ and $4\Rightarrow 1$. Moreover, by [22, Remark 3.2], we know $2\Rightarrow 3$; by [25, Theorem 1.6], we know $3\Rightarrow 4$. Proposition B.2 gives the equivalence between $1$ and $2$. We conclude that the four properties are equivalent. $\Box$